∫cos 2θ log(cosθ+sinθ/cosθ-sinθ)=?

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear ,

 

cosθ + sinθ = √(1 + sin2θ)

cosθ - sinθ = √(1 - sin2θ)

now inregral becomes,  ∫cos2θ*log√(1 + sin2θ) - ∫cos2θlog√(1 - sin2θ)

put √(1 + sin2θ) = t so, cos2θ dθ = tdt

so, ∫cos2θ*log√(1 + sin2θ) = ∫t*logtdt   ...........1

and,

put √(1 - sin2θ) = X so, -cos2θ dθ = XdX

so, ∫cos2θ*log√(1 + sin2θ) = -∫X*logXdX   ...........2

now solve 1 and 2 by parts...

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

vikas askiitian expert
13 years ago

I=cos2@log(cos@+sin@/cos@-sin@) d@

 cos@+sin@ can be written as sqrt{(cos@+sin@)^2}=sqrt(1+sin2@)

 cos@-sin@ can be written as sqrt{(cos@-sin@)^2}=sqrt(1-sin2@)

now I becomes integral cos2@logsqrt{(1+sin2@)/sqrt(1-sin2@)}d@                or

             I=COS2@/2 .LOG(1+SIN2@/1-SIN2@)d@

 now put sin2@=t

     so 2cos2@d@=dt

  I=1/4 .log(1+t/1-t)dt

now using integration by parts

I=1/4{  tlog(1+t/1-t) - integral(1-t)t/1+t dt } 

  =1/4 { tlog1+t/1-t  - integral(2-t-1/t+1)dt

  =1/4{  t.log1+t/1-t  - integral( 2 - t  -2/1+t )dt

  =1/4{ tlog1+t/1-t   -  (2t -t^2/2 -2logt+1)  +c     or

 I=1/4 { sin2@2@(log1+sin2@/1-sin2@)  - 2sin2@ +(sin2@)^2 /2 +2log(1+sin2@)  +c

 approve my ans if u like